OpenStudy (anonymous):

Find the values of the six trigonometric functions of an angle in standard position if the point with coordinates (6,-5) lies on its terminal side.

OpenStudy (anonymous):

In the angle triangle height is -5 and base is 6, So hypotenuse is \[\sqrt{11}\] sin= -5/\[\sqrt{11}\] cos= 6/ \[\sqrt{11}\] tan= -5/6 csc= \[\sqrt{11}\]/8 sec= \[\sqrt{11}\]/6 cot=6/-5

OpenStudy (anonymous):